QUESTION 16 Using the compound interest Formula A= f*1 +5). Find A when P= $ 1000, r=6%, t = 5 years and n= 4 tonas an
Answer #1

P=1000 , r =6% . t = 5 years, n=4 4XS ..A= 1000 (1 + 61 ) = 1000 (1.015) 20 = 1000(1.346855) = 1346.855 7 $ 1346.86